OpenStudy (anonymous):

. What is the distance between the points (10, 3) and (10, -9) in the xy-plane? 12 units 6 units 19 units 10 units

pooja195 (pooja195):

Hey there! Welcome to openstudy! consider reading the CoC: http://openstudy.com/code-of-conduct Heres a link that will help you get to know the site better: http://openstudy.com/study#/updates/548c04e1e4b0f2fbdc23dd6c

OpenStudy (anonymous):

umm..Thank you but I just need help on my question, do you think you can help me?

pooja195 (pooja195):

yes :) which answer choices can you eliminate?

OpenStudy (anonymous):

er.. 6?

jimthompson5910 (jim_thompson5910):

if the x coordinates are the same, then you can simply subtract the y coordinates and take the absolute value to ensure it's positive.

OpenStudy (anonymous):

Ok

pooja195 (pooja195):

ok subtract 10-10 -X values 3-9-Y values

OpenStudy (anonymous):

12?????

pooja195 (pooja195):

no try again however you would get a negative answer but its not here :(

OpenStudy (anonymous):

umm i am in study island and they said 12 was the answer

pooja195 (pooja195):

@jim_thompson5910 can you explain this im confused

jimthompson5910 (jim_thompson5910):

pooja195, distance is never negative

jimthompson5910 (jim_thompson5910):

subtract the y coords and take the absolute value |3 - (-9)| = |3 + 9| = |12| = 12 or |-9 - 3| = |-12| = 12

OpenStudy (anonymous):

Thank you! for the explanition

pooja195 (pooja195):

ah! that explains it thank you :D

OpenStudy (anonymous):

:)

jimthompson5910 (jim_thompson5910):

again, this only works because the x coordinates are the same

pooja195 (pooja195):

right :)